Circular No. 6993

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998dm IN MCG –01-4-44 M. Modjaz, T. Shefler, E. Halderson, J. Y. King, W. D. Li, R. R.

Treffers, and A. V. Filippenko, University of California at Berkeley, report their discovery of an apparent supernova during the course of the Lick Observatory Supernova Search (cf. IAUC 6627) with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). SN 1998dm was found and confirmed on unfiltered images taken on Aug. 22.5 (mag ∼ 16.8) and 23.4 UT (mag ∼ 16.5). The new object is located at α = 1^{h}26^{m}13^{s}.97, δ = −6^{o}06′14′′.0 (equinox 2000.0), which is ∼ 13′′.8 west and 37′′.0 south of the nucleus of MCG –01-4-44 (= UGCA 17). A KAIT image obtained on Aug. 17.5 shows no star at this position (limiting mag ∼ 18.5).

XTE J2123–058 C. Zurita and J. Casares, Instituto de Astrofísica de Canarias (IAC);

and R. I. Hynes, University of Sussex, communicate: “New R-band images of the x-ray transient XTE J2123–058 (IAUC 6955) have been obtained on Aug. 2–16 with the 0.8-m IAC80 telescope at the Observatorio del Teide. Following the behavior of the XTE/ASM lightcurve, the system declined optically from the beginning of August at ∼ 0.1 mag/day. The average magnitude on Aug. 16 was R = 19.1. As noted by Tomsick et al. (IAUC 6957), the object is visible on a digitized U.K. Schmidt plate. We estimate a photographic magnitude fainter than 20 from this image, indicating that the object is likely still several magnitudes above its quiescent state. Despite the drop in luminosity, we observe a dramatic increase in orbital modulation, the current amplitude being 1.4 mag. Two full cycles were recorded on Aug. 15 and 16, showing an almost sinusoidal modulation, consistent with x-ray heating of the companion star. We also note two narrow 0.2-mag dips superposed at phases 0 and 0.5, which could be associated with mutual eclipses between the companion star and accretion disc. The revised ephemeris is HJD 2451042.639(5) + 0.24821(3)E. More observations at other wavelengths are urgently required to follow the evolution of the lightcurve of this unusual transient.”

COMET C/1998 M2 (LINEAR) Total visual magnitude estimates: July 22.95 UT, 13.8 (R. J. Bouma,

Lattrop, The Netherlands, 0.40-m reflector); Aug. 10.87, 14.2 (M. Lehký, Hradec Králové, Czech Republic, 0.42-m reflector); 19.99, 12.8 (J. Carvajal, Toledo, Spain, 0.32-m reflector); 22.93, 12.5 (Carvajal, Avila, Spain).

1998 August 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green